OpenStudy (anonymous):

DEtermine of the given is an arthmetic sequence, if it is find the common difference. n1,2,3,4, f(n) 9,17,25,33 Is it artmetric sequence and what is the common difference.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

17 - 9 = 8 25 - 17 = 8 33 - 25 = 8 @ayushprasad @cdelomas